### v1

# Explanation Bias is a Product: Revealing the Hidden Lexical and Position Preferences in Post-Hoc Feature Attribution

Abstract
Good quality explanations strengthen the understanding of language models and data. Feature attribution methods, such as Integrated Gradient, are a type of post-hoc explainer that can provide token-level insights. However, explanations on the same input may vary greatly due to underlying biases of different methods. Users may be aware of this issue and mistrust their utility, while unaware users may trust them inadequately. In this work, we delve beyond the superficial inconsistencies between attribution methods, structuring their biases through a model- and method-agnostic framework of three evaluation metrics. We systematically assess both the lexical and position bias (what and where in the input) for two transformers; first, in a controlled, pseudo-random classification task on artificial data; then, in a semi-controlled causal relation detection task on natural data. We find that lexical and position biases are structurally unbalanced in our model comparison, with models that score high on one type score low on the other. We also find signs that methods producing anomalous explanations are more likely to be biased themselves.
Explanation Bias is a Product: Revealing the Hidden Lexical and
Position Preferences in Post-Hoc Feature Attribution
Jonathan Kamp1    Roos Bakker2,3    Dominique Blok2 1Computational Linguistics and Text Mining Lab, Vrije Universiteit Amsterdam 2TNO–The Netherlands Organization for Applied Scientific Research, The Hague 3Leiden University Centre for Linguistics (LUCL), Leiden University, Leiden j.b.kamp@vu.nl, {roos.bakker, dominique.blok}@tno.nl
1 Introduction
Figure 1: Bias distributions of top-1 token explanations from ModernBERT classifiers over positions (left; 0–19) and lexical elements (right; "," and ".") in the input. Here we compare two feature attribution methods: Vanilla Gradient (top); Integrated Gradient (bottom). The dotted line indicates the position where we would expect the predictions to be in the absence of bias.
Feature attribution is a post-hoc explanation technique that aims to determine the input tokens that contributed the most to a language model’s prediction. Different attribution methods have been proposed, but they tend to exhibit inconsistent token explanations even on the same input and given the same backbone model (neely2022song). A major problem that comes with attribution inconsistency is the lack of adequate domain knowledge: methods are either adopted without considering the risks of unfaithful explanations (i.e. that do not reflect the true decision mechanisms of the model), or are coarsely deemed unreliable and discarded altogether. Their inconsistent explanation behaviour can manifest itself through various symptoms that reflect some underlying systematic deviation, i.e. a bias, that resides in or surfaces through the attribution method. For example, lyu2024towards suggests that, when methods attribute importance to words such as determiners (the) over adjectives (great), the model is either relying on function words, or the method is not faithfully reflecting the model’s reliance on content words. Since this type of bias concerns lexical information, we refer to it as lexical bias.
However, lexical information might only be a part of the full story, as positional information likely plays a role in bias, too. Figure 1 illustrates the bias distributions for token position and token type of two explanation methods on the same classification task on artificial sequences of periods and commas. Vanilla Gradient (simonyan2014deep) is very biased towards the start of the input texts (index 0). In contrast, the more elaborate Integrated Gradient (sundararajan2017axiomatic) seems more balanced towards position but shows some lexical bias by preferring "." over ",".
In this paper, we study the sources of bias. We argue that lexical and position biases are a product of the attribution method, the model type and the random seed. In this light, we propose an evaluation framework to investigate the biases exhibited by six different attribution methods combined with a traditional and a recent BERT-based classifier. We assess how these transformers, which differ in pre-training, architectural properties and size, suffer from bias: is the modern BERT version less affected? And to what extent are position and lexical bias related? By systematically structuring bias symptoms, we aim to better understand the underlying causes of method–method inconsistencies (also known as disagreement) to help users of attribution methods make more informed decisions. We first experiment in a controlled setting with artificial data, followed by a semi-controlled setting on English causal relation detection.
Among our main findings, we show that these model attributions indeed systematically exhibit certain types of bias. Importantly, a more modern model does not necessarily entail lower bias, and the proneness for a model to exhibit position bias appears to be unrelated to lexical bias, with scenarios in which one increases, the other decreases. Shared results by both artificial and natural settings show that the attribution method producing the most diverging explanations is also the most biased, and that position bias mostly arises towards the end of the inputs in the traditional BERT model. [^1] We publicly release our code at
https://anonymous.4open.science/r/poslex-bias/.
2 Related Work
Attribution methods are inconsistent compared to one another (krishna2022disagreement; neely2022song; kamp-etal-2023-dynamic), which makes it challenging to interpret a backbone model. Possible underlying biases, such as the preference for certain lexical classes and their position in the input, may not only affect model understanding but also the cognitive processing of the presented explanations. For example, jacovi-etal-2023-neighboring carry out a user study showing that lexical and positional information of highlighted tokens are indicative of the way humans value explanations as important.
Lexical bias in attribution methods has been explored in a handful of tasks, such as sentiment analysis and deception detection (lai2019many), and reading comprehension-based question answering (ramnath2020towards). kamp-etal-2024-role investigate the disagreement among attribution methods by mapping explanations to lexical classes. They find, on a natural language inference task, that methods may have different biases towards e.g. nouns, determiners and punctuation but still converge to the same syntactic phrase. They speculate that position bias might be induced by the proximity to an important token. The problem is related to the work of ko-etal-2020-look; ben2024impact, who show that the position of signal words in the input induces a bias towards a specific target class. xiao2023efficient find strong attention scores (sinks), on initial tokens LLM inputs, irrespective of their semantics. eshuijs-etal-2025-short instead find that in a sentiment analysis task, the first most informative word in a signal span receives the most attention in decoder-only models, hence overruling the weight of the word’s position in the span.
The fact that biases are observed on specific tasks raises the question to what extent biases are task- and data-dependent; some contrasting findings by attanasio2022benchmarking versus atanasova2020diagnostic seemed to indicate a task-dependent component in the source of disagreement. We differ from previous work in that we assess both the lexical and position bias in a controlled setting to isolate the findings at best from task-related aspects. To our knowledge, we are the first to do so.
A factor that may directly affect feature attribution is the complexity or size of the model. zhou-etal-2024-explaining explore the effect of training data size, showing that more data can lead to greater plausibility (i.e. alignment with human rationales) but not to greater faithfulness (i.e. alignment with model internals). Notably, they also find that a larger model (Vicuna, 13B (vicuna2023) vs. RoBERTa-large, 355M (liu2019roberta)) does not necessarily produce more plausible or faithful explanations. Conversely, heyen2024the compare LIME (ribeiro-etal-2016-trust) on four versions of DeBERTaV3 (hedebertav3) (22–304M parameters), and find that explanations do not become more plausible in larger models, but do become more faithful. The caveat is that metrics to assess faithfulness can be unreliable, too. in the same study, heyen2024the find that comprehensiveness, which approximates faithfulness as the prediction drop after removing explanation tokens from the input, is not in line with other metrics capturing greater faithfulness for the larger model; this suggests a shortcoming of this metric in specific cases.
We experiment with two BERT-based models. In the process, we also validate two faithfulness metrics: sufficiency and comprehensiveness. The design of the experiments helps us to explore bias in different data constructions with different levels of control; we begin with an artificial setting in §3, followed by a causal relation detection task in §4.
3 Experiments on Artificial Data
We study attribution bias as the frequency in which token explanations are located at certain indices of the input (position bias) and belong to a certain word type (lexical bias). We craft the tasks in a controlled setting in such a way that this information should not play a role in solving the task (which should not be solvable). If attributions are biased, they should manifest this bias consistently across runs and datasets. In §4, we complement this randomised setting with a natural language task.
3.1 Data
We craft three artificial datasets for a binary classification task. For each, we generate a set of 8,000 instances for training, 1,000 for validation and 1,000 for testing. Each instance of fixed length 20 (an average sentence length) is populated with different lexical elements. Within our framework, a set of lexical elements has been chosen that can easily be replaced. The idea is that position bias arises irrespective of the lexical information. Below, we describe these three artificial datasets:
noun-det-period
The chosen lexical classes follow kamp-etal-2024-role, whose initial insights show preferences for specific parts-of-speech. We include a frequent noun (’table’) and determiner (’the’), and a period (’.’). Each token can occur 0+ times. Token positions in the sentences are allocated randomly. Class labels are random and equally distributed.
period-comma
Each instance is a random sequence of ’.’ and ’,’ tokens, each of which can occur 0+ times. The absence of both content and function words may highlight the lexical bias in certain attribution methods that favour punctuation. Class labels are random and equally distributed.
unique-punctuation
Each instance is a unique, random ordering of the set of unique punctuation marks {".", ",", ";", ":", "!", "?", "-", "_", "(", ")", "[", "]", "", "", "/", "*", "#", "’", ’"’, "‘"}. The uniqueness of these lexical elements should facilitate faithfulness assessment (introduced in §3.4) by enabling us to measure the weight of a specific token in the input. Class labels are random and equally distributed.
3.2 Models
We compare two transformer models with a sequence classification head, namely BERT: bert-base-uncased, 110M (devlin-etal-2019-bert) and ModernBERT: ModernBERT-base, 149M (modernbert). We question whether a more recent BERT architecture suffers less from bias. After introducing the bias metrics in §3.4, we speculate further on how model properties may relate to bias.
3.3 Explanations
We compute token attribution scores for each of the input sequences. We ignore start-of-sequence and end-of-sequence tokens, such as [CLS] and [SEP], after computing the attributions in order to focus the analysis on the original 20 tokens per input. We test six different attribution methods: Partition SHAP, or PartSHAP (lundberg2017unified) LIME (ribeiro-etal-2016-trust); Vanilla Gradient, or VanGrad (simonyan2014deep), and Integrated Gradient, or IntGrad (sundararajan2017axiomatic); the latter two multiplied by the input – Grad × I and IntGrad × I shrikumar2017learning. We select the top-$k$ attributions per instance and set $k=1$, i.e. the single strongest attribution preference. This choice ensures a fair comparison with the real world task in §4, where we expect a single word signal to be indicative of the class label.
3.4 Bias Metrics
Given an attribution method, the instance-wise top-1 explanations computed on a test dataset form a frequency distribution over the lexical categories and position indices (Figure 1). These distributions may reflect a bias from the method itself. However, possible biases from pre-training and random weight initialisation would also surface in this way. To properly differentiate between bias types, we propose an analysis framework that includes three key metrics for bias quantification: (a) Bias-cons, the consistency of the frequency bias between the different random seeded runs; (b) Bias-agg, the bias of the aggregate distribution (of a specific attribution method) over 10 fine-tuned models trained with different random seeds; (c) Bias-attr, the bias that is specific to the attribution method, compared to other methods. In principle, the framework is applicable to any attribution method and any top-$k$ value. The proposed baseline is valid in a controlled setting and might need adjustment based on the nature of the task. We use a different baseline for causal relation detection in §4.
Bias-cons for inter-seed comparison
We question to what extent biases are consistent across different random seeds. Since we are interested in the distribution of top-$k$ selections over input tokens, we measure the mean pairwise Jensen-Shannon distance (Bias-cons) between the normalised distributions yielded by different runs. For this, we consider the unique model pairs $(m^{\prime},m^{\prime\prime})$ formed from the set of different runs $\mathcal{M}$ and we average the JS distance over the number of combinations, i.e. $\binom{|\mathcal{M}|}{2}=\frac{|\mathcal{M}|(|\mathcal{M}|-1)}{2}$. JS ranges from 0 (least divergent/biased) to 1 (most divergent/biased). The greater the distance, the greater the impact of random initialisation on the explanations. Formally, for each model type $m\in\{\text{{BERT, ModernBERT}}\}$ and each attribution method $a\in\mathcal{A}$, let $P^{(m^{\prime})}_{m,a}$ or $P^{(m^{\prime\prime})}_{m,a}$ denote the bias distribution (over positions or lexical elements in the input) for a single run $m^{\prime}$ or $m^{\prime\prime}$ (i.e., a specific random seed):
$\text{Bias-cons}_{m,a}=$ (1)
$\frac{1}{\frac{|\mathcal{M}|(|\mathcal{M}|-1)}{2}}\sum_{\begin{subarray}{c}m^{\prime},\,m^{\prime\prime}\in\mathcal{M}\\ m^{\prime}<m^{\prime\prime}\end{subarray}}\text{JS}\left(P^{(m^{\prime})}_{m,a}\,\|\,P^{(m^{\prime\prime})}_{m,a}\right)$
Bias-agg for inter-model comparison
We question the extent to which attribution methods are biased in the light of a model-to-model comparison. For each model type $m$, we compute the Jensen-Shannon distance (JS) between each method $a$’s aggregate bias distribution $P_{m,a}$ and a baseline. $P_{m,a}$ is computed by summing token-level bias distributions across all runs $\mathcal{M}$ and normalising to form a probability distribution. The baseline consists of a normalised uniform distribution $U$ over the positions and lexical elements for the specific dataset, and implies the absence of bias. We refer to this distance metric as Bias-agg, the bias of the aggregated distributions. Aggregate also refers to the fact that this bias measures the conflation of both model bias and attribution method bias.
$\text{Bias-agg}_{m,a}=\text{JS}\left(P_{m,a}\,\|\,U\right)$ (2)
Bias-attr for inter-method comparison
If pre-training were the only source of bias, we would expect all attribution methods to exhibit a consistent bias for the same model, same random seed. This is not the case (attributions are different per method), indicating that at least one or more attribution methods are biased independently from model bias. We define attribution bias (Bias-attr) as the averaged pairwise JS distance between the target method $a$ and each of the other methods $a^{\prime}$ sampled from $\mathcal{A}$.
$\text{Bias-attr}_{m,a}=\frac{1}{|\mathcal{A}\!\setminus\!\{a\}|}\sum_{a^{\prime}\in\mathcal{A}\setminus\{a\}}\text{JS}(P_{m,a}\|P_{m,a^{\prime}})$ (3)
It is unknown which of the attribution methods in $\mathcal{A}$ are the most biased. Based on the principle of reproducibility in science (popper2005logic), a finding that is confirmed through different means is more likely to be correct: can a majority consensus among methods be indicative of faithfulness? Are outlier explanations more likely to be wrong? We therefore question whether we observe similar patterns in Bias-attr and Bias-agg. In other words, are attribution methods that differ mostly from the rest (high Bias-attr) or are most similar to the rest (low Bias-attr) respectively the most and least biased with respect to a baseline expressing absence of bias (high and low Bias-agg)?
Faithfulness metrics as sanity check
Determining the faithfulness of attributions helps us understand whether the explanations reflect the real feature contributions to the model’s prediction. In our controlled setting, a method with high Bias-agg should also be relatively unfaithful, as the model should not have associated any class signals to positions or lexical elements. Do we observe this?
We assess faithfulness through two widely used metrics: sufficiency (suff) (lei2016rationalizing), which is obtained by removing the top-1 tokens (the rationale $r_{i}$) from the input $x_{i}$ and observing the change in probability towards the ground truth label; comprehensiveness (cmp) (yu2019rethinking), which is the change in probability obtained by removing the context ($x_{i}\setminus r_{i}$) around the top-1 tokens from $x_{i}$. Formally, following deyoung2020eraser:
$\text{suff}(D)=\frac{1}{n}\sum_{i=1}^{n}\left(m^{\prime}(x_{i})_{j}-m^{\prime}(r_{i})_{j}\right)$ (4)
$\text{cmp}(D)=\frac{1}{n}\sum_{i=1}^{n}\left(m^{\prime}(x_{i})_{j}-m^{\prime}(x_{i}\setminus r_{i})_{j}\right)$ (5)
where scores are averaged over dataset $D$ and where $m^{\prime}(\cdot)_{j}$ returns a probability by the fine-tuned model $m^{\prime}$ for class $j$ on the full instance ($x_{i}$) and on the corrupted instance ($r_{i}$ or $x_{i}\setminus r_{i}$).
This sanity check assesses the utility of these metrics on poorly performing models, currently an underresearched scenario. We expect the detection of feature faithfulness to be challenging when the model itself does not rely on specific signals.
How will model type relate to bias?
Most of the time, expecting a model to do better merely based on its size is not enough. To speculate more precisely how our models may relate to bias, we take into account the following three considerations. (i) ben2024impact directly relate a form of position bias to the type of positional embeddings in the model. Contrary to BERT, which learnt fixed positional embeddings, ModernBERT uses Rotary Positional Embeddings (RoPE): along with the larger model size, this may lead to a reduced position bias. (ii) Lexical bias may also be reduced by RoPE in ModernBERT, giving more weight to the relative position of a lexical element in the context of the sentence, hence affecting its meaning. (iii) The effect on lexical bias caused by the greater diversity of datasets used by ModernBERT in pre-training is not univocal: it may either mitigate or enforce it.
3.5 Results
As expected, our fine-tuned models achieve chance-level f1-scores, in the range .33–.55. We report the results for all bias metrics based on the top-1 attributions across random seeds for each model type and attribution method. We present the full visuals of the bias distributions in Appendix A.4.
Bias-cons
Results are provided in Figure 2(a) (position bias) and Figure 2(b) (lexical bias). Overall, position bias appears to be more consistent for ModernBERT, displaying a lower Bias-cons in 12 out of 18 scenarios (3 datasets $\times$ 6 methods) and on average (.21 versus .26). The explanations for period-comma are particularly inconsistent for both model types. Conversely, the lexical bias is most consistent for BERT in 11/18 cases and on average (.19 versus .23), with the most visible difference on unique-punctuation.
(a) Position Bias-cons.
(b) Lexical Bias-cons.
Figure 2: Bias-cons: inter-seed consistency on 10 different runs. Low values (light BG) = greater consistency.
Method noun-det-p p-comma uniq-punct noun-det-p p-comma uniq-punct
B ModB B ModB B ModB B ModB B ModB B ModB
Position Bias
PartSHAP .19 .22 .26 .26 .23 .10 .13 .15 .21 .26 .13 .09*
LIME .15 .12 .23 .07 .24 .03 .13 .11 .20 .31 .11 .06
VanGrad .39 .24 .53 .66 .44 .05 .27* .19* .38* .51* .23* .08
Grad × I .18 .11 .37 .36 .30 .05 .13 .11 .23 .26 .15 .08
IntGrad .15 .07 .11 .16 23 .04 .16 .11 .28 .26 .13 .06
IntGrad × I .17 .11 .31 .34 .31 .09 .15 .11 .26 .27 .14 .08
avg. .21 .15 .31 .31 .30 .06
Lexical Bias
PartSHAP .09 .03 .02 .05 .19 .22 .11 .20 .02 .05 .19 .30
LIME .03 .05 .03 .03 .20 .23 .06 .16 .03 .03 .19 .32
VanGrad .04 .30 .06 .01 .34 .56 .07 .26* .02 .07 .31* .54*
Grad × I .02 .16 .06 .04 .16 .34 .06 .14 .02 .09 .20 .36
IntGrad .03 .14 .02 .10 .07 .18 .07 .13 .03 .07 .19 .28
IntGrad × I .12 .21 .04 .11 .19 .21 .14* .17 .07 .08 .26 .30
avg. .05 .15 .04 .06 .19 .29
Table 1: Position and lexical bias comparison between BERT (B) and ModernBERT (ModB) across datasets.
Left, Bias-agg
: per dataset, dark background indicates the model with highest bias (row-wise), including model avg.
Right, Bias-attr
: per dataset, dark background indicates the method with the highest bias (column-wise). (*) indicates high Bias-attr methods (right) that correspond to high Bias-agg (left) for the same model–dataset column.
Bias-agg
Table 1 shows that ModernBERT has the lowest position bias on the three datasets, in 13/18 cases and on average (.15, .31–tie, .06). Notably, proneness to position bias does not seem to be related to lexical bias: similarly to Bias-cons, the roles are reversed, with BERT exhibiting lowest lexical bias in 15/18 cases and on average (.05, .04, .19). The contrast is clear-cut in noun-det-period (including a flipped PartSHAP), mixed in period-comma and clear-cut in unique-punctuation. Despite prior research presenting attribution methods as disagreeing, we find that they show consistency in picking the models with the least and most bias.
Bias-attr
The main pattern in Table 1 is VanGrad having the highest position Bias-attr in 5/6 scenarios, across 3 datasets $\times$ 2 models. VanGrad’s lexical bias is less prominent, but still highest in 3/6 scenarios. There is no clear pattern of lowest bias.
We compare Bias-attr results to Bias-agg results to determine whether the difference between methods is indicative of a difference between the methods and a no-bias-baseline. For this, we consider the highest Bias-attr method per dataset–model pair: in 6/6 cases it corresponds to the highest Bias-agg method for position and in 4/6 cases for lexical bias. If we look at low Bias-attr scores, this parallel pattern with Bias-agg pattern is less clear: several methods exhibit comparably low scores. This indicates that, at least in our experiments, methods with a stronger bias compared to one another (Bias-attr) are also more biased overall compared to a uniform baseline (Bias-agg).
Sufficiency and comprehensiveness
As a sanity check for the faithfulness metrics suff and cmp, we assess whether they are able to detect VanGrad’s high Bias-agg. Table 2 shows that these metrics do not reflect VanGrad having the strongest bias: while suff = .00 would usually indicate faithful explanations, it returns .00 for other methods as well, meaning that no peculiar behaviour of VanGrad is detectable. cmp measures the impact of the context around the top-1 explanation, and understandably is .00 as well, being equally unable to predict the random class as the full input sequence. This confirms our expectations: suff and cmp are not suitable to detect lexical and position bias in attribution methods for models that failed to learn the task (well). By scoring $\approx 0$, these metrics could instead be used as sanity checks for the model, as they correctly reflect that no data artifacts (absent by design) were exploited for the prediction.
B ModB
Method suff cmp suff cmp
PartSHAP -.03 .01 -.07 .01
LIME -.02 .01 -.02 .01
VanGrad -.00 .00 .00 -.00
Grad × I .00 .00 -.04 .01
IntGrad -.01 .00 -.01 .00
IntGrad × I -.00 .00 -.01 .00
Table 2: Sufficiency and comprehensiveness (suff, cmp) for BERT (B) and ModernBERT (ModB) on unique-punctuation. Multiple run averages.
4 Experiments on Causal Relation Data
In Section 3, we analysed position and lexical bias on artificial data in a fully controlled experiment. While a randomised setting provides valuable insights into the sensitivity of attribution methods to exhibit bias under controlled conditions, such findings may not fully capture their behavior in real world tasks. For this reason, we complement this analysis with a semi-controlled experiment on natural language. Specifically, we fine-tune a binary classifier to detect causal relations in a text, using the same BERT and ModernBERT transformers.
Causal relations can be expressed in many ways, both explicitly and implicitly solstadbott. Here we use data with explicit markers, such as causal verbs (cause, depend on) and adverbial phrases (because, as a result of). These strong markers make this task particularly useful for our purpose, contrasting with the artificial dataset where there are no overt markers the models can learn from. There are two main tasks in the domain of causality detection: sentence classification (does it contain causal information?) and causal span detection (cause, marker, and effect) (asghar2016, e.g.). We address the former task.
Position and lexical information are more intertwined in natural language tasks, with their combination forming the semantics that models can encode. One of the main differences is that we cannot expect a uniform distribution as our bias baseline: token types are not evenly distributed and class signals do not appear evenly across indices in the input. We therefore craft the task in such a way that we can study position bias, but focus on sentence position (in a text) rather than token position (in a sentence). This is based on the scenario in which multiple causal signals occur in the same input: which one is going to be responsible for the prediction? After assessing bias in both artificial and natural language experiments, we carefully compare whether any bias patterns hold for the same models and attribution methods.
4.1 Data
The data is adapted from veldhuisetal, and consists of annotated texts that contain causal relations used for system dynamics modelling. They stem from documents on climate change, putinism policy and traffic congestion (sterman2003climatechange; mcfaul2020putin; sterman2000trafficcongestion). We finetune the classifiers on alterations of these texts to i) emulate the repetition of causal signals in the same input and ii) to control for position bias compared to a baseline. For this, we first randomly sample sentence triples (900 for training, 100 for testing). We then construct input instances as permutations of these triples, with each triple resulting in 6 different orderings of the three different sentences (5400 for training, 600 for testing). Positive instances (50% of the data) consist of concatenations of three different positive sentences (+++); the same logic applies to negative instances (---). The three sentences in +++ represent causal occurrences at beginning-, middle- and end-of-text. We name this diagnostic dataset causal. The task is to predict a positive or negative class label for each instance. We train and test on the perturbations so that the model should not learn that the dominant causal signals (each instance is expected to contain one) are tied to specific sentence positions. By only changing the order of the sentences (and not e.g. the syntactic structure) and keeping the tokens constant per permutation batch, we control for confounding factors that would arise from feeding different inputs. On average, positive sentences contain 24.6 words (6–60), negative 18.4 (1–57), while +++ and --- test instances contain 64.1 (22–128). Example inputs are given in Appendix A.1.
4.2 Bias and Baseline Distributions
Position bias is measured as a distribution over the three sentence positions in each of the inputs. In positive instances, each sentence will contain a signal of causality. It is reasonable to expect that the causal signal of one particular sentence will be dominant. For each set of 6 permutations, the dominant signal would appear twice in the sentence placed first, twice second and twice third. Therefore, the baseline implying absence of position bias is a uniform distribution over sentence positions.
To deepen our analysis, we differentiate between positive class attributions and negative class attributions. The former are expected to target strong causal signals, whereas the latter must deal with a signal absence, which makes attribution intuitively more challenging (e.g. dela2024none).
4.3 Results
BERT and ModernBERT fine-tuned on the causal task achieve f1-scores $\scriptstyle>$.98, hence forming an interesting counter scenario to the artificial experiments (f1-score $\scriptstyle\sim$.50). From a first lexical insight, we notice that targeted top-1 tokens vary across models and methods. For example, BERT’s LIME selects mostly causal verbs such as produce, increases, achieve, influence, depends, weaken, triggered. ModernBERT’s IntGrad selects a mix of different lexical elements, such as causal verbs (made, change, developed), but also punctuation, determiners and topic-related words (temperature, climate, regime, emissions). The latter implausible explanations increase the relevance of investigating where in the input they occur. Moreover, methods attribute more function words and punctuation in negative examples, making up for the absence of causal signals. In this section, we report the sentence position biases from these models.
(a) Positive class Bias-cons for sentence position.
(b) Negative class Bias-cons for sentence position.
Figure 3: Bias-cons: inter-seed consistency on 10 different runs. Low values (light BG) = greater consistency.
(a) BERT, positive class.
(b) ModernBERT, positive class.
(c) BERT, negative class.
(d) ModernBERT, negative class.
Figure 4: Sentence position Bias-agg (relative to input triple) in causal relation detection. Top row: positive class (+++, causal); bottom row: negative class (---, non-causal). Vertical red line indicates chance.
Bias-cons
The scores in Figure 3 are generally low (.03–.15 across models): attribution methods yield a relatively consistent bias across sentence positions. This means that when a task is learnt successfully, a dominant causal signal is most likely identified irrespective of position. Low scores are in part due to a relatively low number of sentence positions (3), facilitating the task by conflating token attributions into sentence attributions. We notice that whereas LIME and IntGrad were roughly the most consistent on the artificial data, only LIME is on causal, suggesting that LIME is only lightly affected by the models’ different random seeds. Instead, the least consistent methods on the artificial sets (VanGrad for BERT, IntGrad × I for ModernBERT) are not the least on causal.
Different from the artificial setting, ModernBERT is not predominantly less biased than BERT, despite BERT’s high bias on IntGrad and IntGrad × I (.13 and.15). Finally, positive class biases are more consistent for BERT, while negative examples are marginally more consistent for ModernBERT.
Bias-agg and Bias-attr
For BERT, the attributions point to the dominant (non-)causal signal appearing with a slight bias in the last sentence of each +++ or --- instance (Figure 4). This suggests that BERT is most biased towards late information in the input, especially on negative examples, which reflects the results on artificial data (Figures 5, 6, 7, Appendix A).
ModernBERT is overall less biased than BERT, which also is in line with the artificial setting (Table 1). This is especially visible on negative examples (Bias-agg .01–.03), with methods being aligned on this bias pattern (Bias-attr .02–.03). Interestingly enough, Bias-agg is higher and more variable on ModernBERT $\times$ positive class, with PartSHAP (.11) and LIME (.01) standing in sharp contrast.
In line with the artificial data, we find a similar relation between Bias-agg and Bias-attr: the method with the highest Bias-attr also exhibits the highest Bias-agg for both models on positive class (IntGrad–Fig. 4(a); PartSHAP–Fig. 4(d)), and on the top two methods on BERT $\times$ negative class (PartSHAP and IntGrad × I–Fig. 4(c)). In Fig. 4(d), a pattern is hardly found, but biases are already low.
5 Conclusion and Future Directions
We proposed a bias evaluation framework on both artificial and natural language data, finding systematic preferences for lexical and positional information. We questioned how lexical and position bias relate and whether a recent BERT transformer would be less biased than its traditional version. We conclude that the two types of biases do not imply one another and that ModernBERT is not necessarily less biased than BERT. We also questioned whether evidence from anomalous explainers is more likely to be incorrect. To this end, we found repeated empirical evidence that methods differing most from other methods are also the most biased. Since our settings controlled for lexical and position confounders, this bias is expected to stem from the method itself and not from the backbone model.
[…]
Limitations
We report some limitations of our work. The first concerns the sample of feature attribution methods that were adopted in our analyses; while it includes methods from both gradient-based and perturbation-based families, they are not directly representative of other methods outside this sample. The risk exists that our results are inadequately generalised for other explainable AI methods, even unrelated to feature attribution. Similarly, we crafted multiple data variations for the artificial setting and observed their common traits as a showcase of our proposed framework; however, the possibilities for non-natural lexical combinations are potentially infinite, and may be refined with more precise hypotheses in future work. Again, the risk exists that the results from these specific datasets are generalised over any controlled setting with random token sequences. Finally, controlled experimental settings favour phenomenon isolation but also limit the degrees of comparison with less controlled settings; this applies to the natural language task, which had to rely on different baseline assumptions (non-uniform lexical preferences) and different language units (sentences versus tokens) than the controlled, artificial setting.
Acknowledgments
This research is funded by the Dutch National Science Organisation (NWO) through project code NWA.1292.19.399 – InDeep: Interpreting Deep Learning Models for Text and Sound. We would also like to thank TNO and the APPL.AI programme for their support of this project.
Appendix A Appendix
A.1 Example Inputs
We provide a positive (+++) and a negative (---) input instance from the causal dataset on which predictions and attributions were computed.
+++ Still more believe that stabilizing emissions near current rates would stabilize the climate, when in fact emissions would continue to exceed removal, increasing GHG concentrations and radiative forcing. The amount of energy radiated back into space depends on the composition of the atmosphere. People tend not to enjoy longer travel times, thus, once travel times increase a pressure will build up to reduce it.
--- Both of these explanations allow no role for the agency of individual leaders and their ideas in the analysis. At the Munich Security Conference in 2016, Prime Minister Dmitry Medvedev referenced the Cuban missile crisis as a similar moment in bilateral tensions: Speaking bluntly, we are rapidly rolling into a period of a new cold war. The new Syrian generals empowered by a negotiated transition would likely have remained loyal to Moscow.
A.2 Hardware and Runtimes
The runtimes for fine-tuning the classifiers and generating the explanations are reported in Table 3. All models were fine-tuned on NVIDIA GeForce RTX 2080 Ti (11264 MiB), which was also used to generate the explanations. We used a batch size of 8 and we set the learning rate to 5e-6 with a decay factor of 1e-2.
Model Size fine-tuning explaining
artificial
BERT 110M 2 min $\sim$2–5 sec
ModernBERT 149M 3 min $\sim$2–5 sec
causal
BERT 110M 6 min $\sim$7–9 sec
ModernBERT 149M 8 min $\sim$16–18 sec
Table 3: Model fine-tuning times (5 epochs) on artificial and causal data along with explanation generation times per instance (6 default methods in Ferret (attanasio-etal-2023-ferret) with default parameters).
A.3 Data Artifacts
We use the existing dataset by veldhuisetal. It was obtained on request to the authors, who agreed to publish the data splits used in our experiments on our repository. We advice to contact the authors for re-use of the data, intended for research purposes. The data was manually checked for offensive content or personal information, which were not encountered.
A.4 Complementary Visualisations
We provide the full overview of bias distributions for all attribution methods and model types in Figure 5, Figure 6 and Figure 7 (one for each artificial dataset). These form the foundation for the bias computations reported in §3.5.
(a) BERT
(b) ModernBERT
Figure 5: noun-det-period biases. Different attribution methods show different top-1 preferences for lexical elements and their position in the sentence. Left: position frequency distribution. Right: lexical frequency distribution. Vertical red line indicates chance.
(a) BERT
(b) ModernBERT
Figure 6: period-comma biases. Different attribution methods show different top-1 preferences for lexical elements and their position in the sentence. Left: position frequency distribution. Right: lexical frequency distribution. Vertical red line indicates chance.
(a) BERT
(b) ModernBERT
Figure 7: unique-punctuation biases. Different attribution methods show different top-1 preferences for lexical elements and their position in the sentence. Left: position frequency distribution. Right: lexical frequency distribution. Vertical red line indicates chance.